### v3

Abstract
Accurately reconstructing a 3D scene including explicit geometry information is both attractive and challenging. Geometry reconstruction can benefit from incorporating differentiable appearance models, such as Neural Radiance Fields and 3D Gaussian Splatting (3DGS). However, existing methods encounter efficiency issues due to indirect geometry learning and the paradigm of separately modeling geometry and surface appearance. In this work, we propose a learnable scene model that incorporates 3DGS with an explicit geometry representation, namely a mesh. Our model learns the mesh and appearance in an end-to-end manner, where we bind 3D Gaussians to the mesh faces and perform differentiable rendering of 3DGS to obtain photometric supervision. The model creates an effective information pathway to supervise the learning of both 3DGS and mesh. Experimental results demonstrate that the learned scene model not only improves efficiency and rendering quality but also enables manipulation via the explicit mesh. In addition, our model has a unique advantage in adapting to scene updates, thanks to the end-to-end learning of both mesh and appearance. The code is available at https://github.com/Cenbylin/DMGS.
keywords:
3D reconstruction, Gaussian splatting, Hybrid representation, Mesh-guided manipulation
1 Introduction
Mesh is often the preferred choice in numerous geometry processing scenarios, as it carry explicit surface information. When building a 3D scene model that includes a mesh, it is challenging to directly learn from visual observation due to the complicated pipeline required to produce the synthetic views. Recently, neural surface reconstruction methods [1, 2, 3] have been developed to represent scene geometry as an implicit signed distance function (SDF). These methods establish a connection with Neural Radiance Fields (NeRFs) [4], enabling effective image-based supervision for geometry learning. However, the geometry is learned indirectly, as it involves converting the SDF to a radiance field. Additionally, NeRF-based learning requires expensive ray-marching in the volumetric rendering. These factors lead to efficiency challenges in learning geometry and appearance.
3D Gaussian Splatting (3DGS) [5] has recently gained popularity due to its ability to render photorealistic images significantly faster than NeRFs. 3DGS employs an explicit representation called Gaussians, which represent anisotropic ellipsoids in 3D space. Gaussians enable efficient rendering due to their explicit nature and the highly parallelized splatting procedure. Despite the advantages in rendering quality and speed, it is not straightforward to reconstruct object surfaces in the traditional form of Gaussians. This is because the 3DGS model imposes no explicit constraints on the structure of the Gaussians. As illustrated in Fig. 1, Gaussians that result in visually appealing rendering may be geometrically inaccurate: they can be inconsistent with any 2D manifold. Recently, some methods have incorporated regularization into the learning process of 3DGS. NeuSG [6] refines Gaussians using surface normals estimated by a jointly trained NeuS model [1], which notably increases the training time. SuGaR [7] encourages Gaussians to align with the surface and subsequently converts them into a point cloud for Poisson Surface Reconstruction. Once a coarse mesh is extracted, an additional stage is required to learn surface appearance from scratch using a new set of Gaussians. This two-stage pipeline may introduce efficiency problems. It is worth noting that similar issues are encountered in existing downstream applications [8, 9, 10, 11], which perform meshing and Gaussian training separately. These challenges motivate us to design a more effective information pathway to supervise the learning of scene representations.
Figure 1: The Gaussians and the underlying surface did not align well in the original 3DGS [5], whereas our hybrid representation explicitly restricts the Gaussians to the mesh faces. Our method benefits from the high-quality rendering of 3DGS and the explicit surface structure provided by a mesh.
In this paper, we present a technique to explicitly bind Gaussians to mesh faces, enabling their simultaneous learning. Specifically, the mesh is extracted from a learnable SDF of objects using differentiable marching techniques. The 3D Gaussians are then constructed as disks that are regularly attached to the mesh faces, as shown in Fig. 1 (b). Since the 3DGS model is bound to the zero-level set of the SDF representing object surfaces, the photometric supervision applied to the 3DGS model simultaneously guides the reconstruction of both the appearance and geometry of a scene. In addition, we model the background elements using the original 3DGS, which is jointly learned with our proposed surface-bound Gaussians. This approach makes our method more practical compared to existing methods that require inconvenient foreground masks [12, 13]. As the mesh faces vary during training, the Gaussians dynamically bound to them struggle to learn per-Gaussian colors as effectively as in the original 3DGS. To address this issue, we employ a neural network to learn the scene’s appearance and predict Gaussian colors for rendering.
Extensive experiments show that our method achieves superior rendering accuracy and produces high-quality meshes in indoor and outdoor scenes. Moreover, it reduces total reconstruction time by 50% compared to recent methods and is over 10x faster than methods that rely on Neural SDFs and NeRFs [14, 1]. Additionally, we show that our method offers a unique advantage in adapting to scene updates, outperforming existing two-stage methods [7, 15].
In summary, our contributions are as follows:
  • We present a framework for directly learning scene representations with mesh and Gaussian-based appearance in an end-to-end manner, while ensuring explicit alignment and correspondence between Gaussians and mesh faces.
  • We propose a technique to model background appearance, extending the applicability of differentiable mesh extraction methods to outdoor scenes without the need for foreground masks.
  • We introduce neural appearance model that predicts Gaussians colors, enhancing the robustness of the optimization process.
2 Related Work
2.1 Scene Reconstruction via Radiance Field
Neural Radiance Field (NeRF) [4] is a significant advancement in 3D reconstruction. It uses neural networks to model the radiance field and yield high-fidelity images via volumetric rendering. Despite its outstanding rendering quality, NeRF requires a long time for training and rendering due to the large number of samples needed to query the neural network. Recent efforts to accelerate NeRF have employed explicit grids [16] or hybrids of grids and small MLPs [17, 18, 19]. However, these computational accelerations might compromise image quality. In addition to improving speed, some works explore scene manipulation and composition within NeRF [20], but these are still not convenient for downstream applications. This has led to another series of works [21, 1] that connect NeRF with a mesh, which will be discussed in the next section.
Recently, 3D Gaussian Splatting (3DGS) [5] employs anisotropic Gaussians to model an explicit radiance field, achieving real-time rendering and enhanced quality. The explicit nature of 3DGS makes it suitable for various domains such as scene understanding [22, 23] and AIGC [24], where efficiency and controllability are crucial. However, directly manipulating 3DGS is challenging due to the vast number of Gaussian components involved. Hence, existing applications like Avatars [8], VR [9], and physics simulations [10, 11] leverage mesh guidance to manipulate the scene more effectively.
2.2 Image-based Surface Reconstruction
Reconstructing surfaces from images is a fundamental task in computer vision. Traditional methods, such as Multi-view Stereo (MVS) [25, 26, 27], estimate depth maps or extract voxel grids based on correspondences between images. However, their quality is heavily dependent on the accuracy of image matching, and the use of voxel grids is often constrained by their cubic memory requirements. Recent advancements [12, 13] utilize differentiable isosurface extraction to directly optimize surface meshes under image supervision. Nonetheless, these methods are mainly for single objects due to the substantial memory demands of larger scenes.
Another promising direction, as demonstrated by NeuS [1, 28, 14], adopts a hybrid representation by transforming a Signed Distance Function (SDF) into a density field, adopting the learning pipeline of NeRFs. These techniques can model view-dependent appearances and also allow for the extraction of meshes from the SDF, making them suitable for downstream applications. To enable real-time rendering, some approaches [2, 29, 21] bake the appearance from a trained NeRF into the mesh textures. However, they typically require extensive training times and might reduce rendering quality.
Several studies have achieved mesh reconstruction using 3DGS models. DreamGaussian [30] evaluates the density value of grid nodes using a mixture of neighborhood Gaussians and employs the Marching Cubes algorithm [31] to extract the isosurface. However, similar to NeRFs, the resulting density field may not accurately reflect the surface, leading to unsatisfactory surface quality in complex scenes. NeuSG [6] jointly learns NeuS and 3DGS models while introducing a regularization term to ensure consistency between the Gaussian orientation and the surface normal predicted by NeuS. This joint optimization significantly extends the training duration to over ten hours. Most recently, SuGaR [7] regularizes Gaussians using derived SDF values to align them with the surface and then treats the Gaussians as a point cloud for Poisson Surface Reconstruction. However, SuGaR requires the training of an additional set of Gaussians that are bound to the extracted mesh faces.
We also highlight the differences between our method and the recent works SuGaR [7] and GaMeS [15], which both construct a hybrid representation of Gaussians and mesh. Existing works first extract the mesh or directly use the ground-truth mesh and then learn Gaussian-based appearance from scratch. In contrast, our approach simultaneously learns such a hybrid representation in an end-to-end manner. We will demonstrate that our method achieves a more efficient pipeline and better performance.
3 Preliminary
3.1 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [5] represents the 3D scene using a collection of Gaussians, each defined as a density function in $\mathbb{R}^{3}$:
$G(\boldsymbol{x})=\exp\left\{-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{T}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})\right\}$ (1)
where $\boldsymbol{\mu}$ is a center position (mean), and $\boldsymbol{\Sigma}$ denotes a 3D covariance matrix. To model scene appearance, each Gaussian is associated with an opacity $\alpha\in[0,1]$, and a view-dependent color $\boldsymbol{c}$, which is parameterized with a set of Spherical Harmonics coefficients (SHs) $\{\boldsymbol{y}_{l}^{m}\in\mathbb{R}^{3}\}_{0\leq l\leq l_{\text{max}}}^{-l\leq m\leq l}$ [16].
The interaction of a ray with a Gaussian can be characterized by integrating the Gaussian’s density function $G(\boldsymbol{x})$. This leads to an efficient imaging model that projects Gaussians into 2D image space, resulting in 2D Gaussians with covariance calculated as:
$\boldsymbol{\Sigma}^{\prime}=\boldsymbol{JW\Sigma}\boldsymbol{W}^{T}\boldsymbol{J}^{T},$ (2)
where $\boldsymbol{W}$ represents the current viewing transformation, and $\boldsymbol{J}$ is the Jacobian for an affine approximation to the projective transformation [32].
To query the color of a pixel at location $\boldsymbol{x}^{\prime}\in\mathbb{R}^{2}$ on image $\mathbf{I}$, a sorted list $\mathcal{S}$ of contributing Gaussians is determined based on their depths. The color of this pixel is then computed as follows:
$\mathbf{I}(\boldsymbol{x}^{\prime})=\sum_{i\in\mathcal{S}}\boldsymbol{c}_{i}\alpha_{i}^{\prime}\prod_{j=1}^{i-1}(1-\alpha_{j}^{\prime}),$ (3)
where $\boldsymbol{c}_{i}$ is the color of the $i$-th Gaussian, and $\alpha_{i}^{\prime}$ is the opacity contributed by the Gaussian, computed by:
$\alpha_{i}^{\prime}=\alpha_{i}\exp\left\{-\frac{1}{2}(\boldsymbol{x}^{\prime}-\boldsymbol{\mu}_{i}^{\prime})^{T}\boldsymbol{\Sigma}_{i}^{\prime-1}(\boldsymbol{x}^{\prime}-\boldsymbol{\mu}_{i}^{\prime})\right\},$ (4)
where $\alpha_{i}$ and $\boldsymbol{\mu}_{i}^{\prime}$ denote the learnable opacity and the projected center of the $i$-th Gaussian, respectively.
Using images from known camera poses, the parameters of the Gaussians are adjusted to align rendered and observed images $\mathbf{I}_{\text{GT}}$ by minimizing the photometric loss:
$\underset{\alpha,\boldsymbol{\mu},\boldsymbol{c},\boldsymbol{\Sigma}}{\text{minimize }}\mathcal{L}_{\text{photo}}(\mathbf{I},\mathbf{I}_{\text{GT}}).$ (5)
The optimization variables for covariance matrix $\boldsymbol{\Sigma}$ include a quaternion $\boldsymbol{q}$ and a 3D vector $\boldsymbol{s}$, which relate to the objective through a change of variables:
$\boldsymbol{\Sigma}=\boldsymbol{R}\boldsymbol{S}\boldsymbol{S}^{T}\boldsymbol{R}^{T},$ (6)
where $\boldsymbol{R}$ and $\boldsymbol{S}$ are the rotation and scaling matrices derived from $\boldsymbol{q}$ and $\boldsymbol{s}$.
3.2 Signed Distance Function
The Signed Distance Function (SDF) is a scalar field that defines the distance from any point in space to the nearest surface of an object. The sign of the distance indicates whether the point is inside or outside the object. Formally, given a point $\boldsymbol{x}$ in 3D space, the SDF value is defined as:
$D(\boldsymbol{x})=\begin{cases}+d(\boldsymbol{x},\partial\Omega)&\text{if }\boldsymbol{x}\in\mathbb{R}^{3}\setminus\Omega\\ -d(\boldsymbol{x},\partial\Omega)&\text{if }\boldsymbol{x}\in\Omega\end{cases}$ (7)
where $d(\boldsymbol{x},\partial\Omega)$ is the Euclidean distance from $\boldsymbol{x}$ to the surface $\partial\Omega$ of the object $\Omega$.
Directly learning a mesh from images is challenging. A common paradigm is to optimize an parameterized SDF and then extract a triangle mesh approximating the level set of that function. Notable examples include recent works that use differentiable marching techniques [33, 13, 12] and the conversion from SDF to neural radiance fields [14, 1, 28].
4 Method
[…]
Figure 2: Method Overview. The mesh is derived from a learnable SDF grid using a differentiable marching algorithm. Gaussians are created from the mesh faces, ensuring their alignment with the surface. A neural appearance model determines colors for Gaussians, which are then used to render an image.
The proposed method aims to obtain a set of Gaussians that strictly align with object surfaces. Specifically, these surfaces are constructed with triangle meshes, and the Gaussians are precisely defined in relation to the mesh faces.
As illustrated in Fig. 2, our method includes modular differentiable steps, allowing the model to learn from images as in the original 3DGS. Let the 3D scene of interest be within a region $U\subset\mathbb{R}^{3}$. The model consists of:
  1. 1.
    a scalar field of signed distance function $D:U\mapsto\mathbb{R}$, which relates locations to the presence of objects;
  2. 2.
    a triangle mesh $G=(\mathcal{V},\mathcal{F})$ derived from $D$, where vertices $\mathcal{V}=\{\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{V}\}\subset\mathbb{R}^{3}$ and faces =$\mathcal{F}=\{\boldsymbol{f}_{1},\dots,\boldsymbol{f}_{F}\}$, with each $\boldsymbol{f}$ denoting a triplet of indices within $[1,\dots,V]$;
  3. 3.
    a constrained 3DGS model where the Gaussians are attached to the faces in $\mathcal{F}$;
  4. 4.
    a neural appearance model, which maps a Gaussian center to spherical harmonics coefficients.
4.1 Signed Distance Function and Differentiable Mesh Computation
The geometry of the scene consists of object surfaces, represented as the zero-level sets of the signed distance function, denoted by $\{\boldsymbol{x}|D(\boldsymbol{x})=0\}$. We represent $D$ using an explicit format: for a set of points $\{\boldsymbol{x}_{n}\}_{n=1}^{N}$, which correspond to the nodes of a regular 3D grid, the function values $\{s_{n}\}_{n=1}^{N}$ are specified. For a generic location $\boldsymbol{x}$, $D(\boldsymbol{x})$ is computed by interpolation from the eight nodes of the grid cell enclosing $\boldsymbol{x}$. Therefore, the value set $\{s_{n}\}_{n=1}^{N}$ forms a parametric representation of $D$.
Given the SDF, the object surfaces are implicitly defined by its zero-level set. However, a practical 3D scene model requires explicit geometry to be incorporated into the computational pipelines for rendering and learning. A traditional method to recover a discrete surface representation for a given SDF is the Marching Cubes algorithm [31], which results in a triangle mesh consistent with a 2D manifold. In contrast, this work involves not just recovering a surface but also learning the SDF. This requires the model to adjust the SDF parameters $\{s_{n}\}_{n=1}^{N}$ to make the resultant surfaces vary according to desired updates. Therefore, we adopt the recently proposed FlexiCubes [13] as the differentiable surface extractor.
Additionally, we address two substantial challenges of modeling outdoor and realistic scenes when learning the SDF: (1) For real-world scenes, especially unbounded outdoor environments, the scale of the background ($\mathbb{R}^{3}\setminus U$) is significantly large and costly to represent using an explicit mesh unified with the foreground $U$. Existing methods often manually exclude the background using known foreground segmentation masks [12, 13]. This approach is inconvenient in practice, and visual modeling of the background is still necessary for tasks such as scene image synthesis. (2) As the scale of $U$ increase, the memory required to execute differentiable marching algorithms grows significantly.
We address challenge (1) by incorporating an additional set of Gaussians to handle the visual components from the background. Specifically, we construct background Gaussians $\mathcal{GS}_{\text{bg}}=\{\alpha_{k},\boldsymbol{\mu}_{k},\boldsymbol{c}_{k},\boldsymbol{\Sigma}_{k}\}$ using the original form of 3DGS. Due to their explicit nature, $\mathcal{GS}_{\text{bg}}$ can be directly combined with the foreground Gaussians $\mathcal{GS}_{\text{fg}}$ attached to the mesh (detailed in the next section), enabling rendering and training.
For challenge (2), we apply two schemes in the use of the SDF grid to mitigate the storage and computational demands, thus enabling a larger scale of $U$:
[…]
Visible Node Optimization During each training iteration, only the scene elements within a single camera’s viewing frustum are learning. Therefore, we extract mesh faces only from the visible grid nodes in $\{\boldsymbol{x}_{n}\}_{n=1}^{N}$ using the camera’s intrinsic and extrinsic parameters.
[…]
Coarse-to-fine Grid Starting with a high-resolution grid can lead to redundant faces, causing unstable learning and memory issues. To address this, we begin with a low-resolution grid and progressively increase the resolution during training. The resolution is enhanced by linearly interpolating between the nodes of the current grid and the refined grid.
More technical details can be found in the appendix.
Figure 3: Sub-figure (a) shows using $K=1,3,6$ Gaussians to represent the appearance of a triangle face, (b) defines a local coordinate frame, and (c) illustrates applying linear transformation $\boldsymbol{M}$ to let Gaussians adapt to the irregular triangle. See Sec. 4.2 for more details.
4.2 Gaussians Consistent with Scene Geometry
As shown in Fig. 3 (a), we employ $K$ Gaussians to represent the surface appearance within a triangle $(\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3})$, where $K$ depends on desirable level of detail.
The centers of the Gaussians are placed at specific barycentric coordinates within the triangle. Let these barycentric coordinates be $\boldsymbol{\xi}_{1},\boldsymbol{\ldots},\boldsymbol{\xi}_{K}\in\mathbf{S}^{2}$, where $\mathbf{S}^{2}$ is the 2-simplex: $\mathbf{S}^{2}=\{(x_{1},x_{2},x_{3})|x_{n}\geq 0,x_{1}+x_{2}+x_{3}=1\}$. In the world coordinate system, the center of a Gaussian is computed as:
$\boldsymbol{\mu}_{k}=[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}]\cdot\boldsymbol{\xi}_{k}.$ (8)
Note that although the barycentric coordinates $\{\boldsymbol{\xi}_{k}\}$ are fixed, the centers $\{\boldsymbol{\mu}_{k}\}$ depend on the mesh vertices $\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}$, which in turn depend on the SDF field. The configurations of $\{\boldsymbol{\xi}_{k}\}$ for different $K$ values can be found in the appendix.
Before determining the Gaussian covariances, we introduce a local coordinate frame as illustrated in Fig. 3 (b). This frame has its origin at $\boldsymbol{v}_{1}$. The axes of the frame are denoted as $\boldsymbol{t}_{i}=\tilde{\boldsymbol{t}}_{i}/\|\tilde{\boldsymbol{t}}_{i}\|$, $i\in\{1,2,3\}$, where
$\tilde{\boldsymbol{t}}_{1}=\boldsymbol{a}\times\boldsymbol{b},\quad\tilde{\boldsymbol{t}}_{2}=\boldsymbol{a},\quad\tilde{\boldsymbol{t}}_{3}=(\boldsymbol{a}\times\boldsymbol{b})\times\boldsymbol{a},$ (9)
where $\boldsymbol{a}\equiv\boldsymbol{v}_{2}-\boldsymbol{v}_{1}$, $\boldsymbol{b}\equiv\boldsymbol{v}_{3}-\boldsymbol{v}_{1}$, and $\times$ denotes the cross product. The rotation matrix from local space to world space is then $\boldsymbol{R}_{t2w}=[\boldsymbol{t}_{1},\boldsymbol{t}_{2},\boldsymbol{t}_{3}]$.
If the angles of a triangle are all close to $\pi/3$, the Gaussian covariances in the local coordinate frame can be naturally defined as $\boldsymbol{\Sigma}_{e}=\text{diag}(\epsilon,r^{2},r^{2})$, representing a thin layer of Gaussian surrounding a small area of the surface. The choice of $r$ is empirical. For triangles with a more irregular aspect ratio, where the area is challenging to approximate by a disc, it is beneficial to apply a linear transformation $\boldsymbol{M}$ on $\boldsymbol{\Sigma}_{e}$ as visualized in Fig. 3 (c). Hence, the covariance matrix in world coordinate system is given by:
$\boldsymbol{\Sigma}=\boldsymbol{R}_{t2w}\boldsymbol{M}\boldsymbol{\Sigma}_{e}\boldsymbol{M}^{T}\boldsymbol{R}_{t2w}^{T}.$ (10)
The derivation of matrix $\boldsymbol{M}$ can be found in the appendix.
4.3 Gaussian Appearance Model
In the original 3DGS model, the per-Gaussian colors $\{\boldsymbol{c}_{k}\}$ are optimized directly because the Gaussians remain consistent throughout the training process. However, in our framework, the Gaussians are dynamically constructed from the evolving SDF. Unlike scene geometry, where gradients from the Gaussian centers $\{\boldsymbol{\mu}_{k}\}$ flow back through the same parametric representation $\{s_{n}\}_{n=1}^{N}$ of the SDF, this dynamic construction makes continuous color optimization difficult.
Figure 4: Comparison between two approaches to learn appearance.
Specifically, the key challenges we face is the lack of direct connections between Gaussians derived from $(\mathcal{V}_{t},\mathcal{F}_{t})$ and those from $(\mathcal{V}_{t-1},\mathcal{F}_{t-1})$, where $t$ denotes the $t$-th training step. A straightforward approach to overcome this would be to compute colors based on their neighboring Gaussians from the previous step, formulated as $\boldsymbol{c}_{i,t}=\frac{1}{|\mathcal{N}|}\sum_{j\in\mathcal{N}}\boldsymbol{c}_{j,t-1}$, where $\boldsymbol{c}_{i,t}$ denotes the color of the $i$-th Gaussian at step $t$, and $\mathcal{N}$ is the set of neighbors based on the Gaussian centers. However, this approach presents two significant drawbacks: it requires substantial computation due to the neighbor search between two large sets of Gaussians, averaging colors within a local region might lead to low effectiveness in learning scene details, as shown in Fig. 4 (b).
To address these challenges, we introduce a neural appearance model, $\mathcal{A}:\mathbb{R}^{3}\mapsto\mathbb{R}^{d}$, with $d$ representing the number of the used Spherical Harmonics coefficients. At each step, we determine the color as $\boldsymbol{c}=\mathcal{A}(\boldsymbol{\mu})$ for the Gaussian centered at $\boldsymbol{\mu}$. Thus, the appearance of Gaussians depends on the learnable parameters within $\mathcal{A}$, which remain consistent during training.
We implement $\mathcal{A}$ using hash-grid positional encoding [18] and a lightweight MLP to enhance efficiency and achieve a compact and continuous representation, inspired by NvdiffRec [12]. This neural appearance model can be robustly learned, even as the mesh faces changes from $(\mathcal{V}_{t-1},\mathcal{F}_{t-1})$ to $(\mathcal{V}_{t},\mathcal{F}_{t})$. Fig. 4 illustrates the experimental comparison between the two approaches. The neural appearance model can achieve nearly 40x speed-up and better learned details than the straightforward nearest neighbor searching approach.
Additionally, we set the opacity to $\alpha=1$ to avoid semi-transparent faces.
4.4 Rendering and Optimization
As the Gaussians we construct retain the same form as those in the original 3DGS, we employ the same method defined in Eq. (3) for rendering an image. For scenes with a background, the rendered Gaussians include both the Gaussians derived from the mesh and the additional background Gaussians. Our method relies only on photometric supervision and utilizes the loss function:
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM},$ (11)
where $\lambda$ is a weighting factor, and $\mathcal{L}_{1}$, $\mathcal{L}_{D-SSIM}$ represent standard L1 and D-SSIM metrics for comparing the rendered image against the ground truth.
Upon the convergence of the mesh and appearance model, we fix the topology defined by faces $\mathcal{F}$ and proceed to jointly refine the vertices and the associated 3D Gaussians. During this phase, Gaussian centers continue to be derived from the face vertices using established barycentric coordinates. However, for covariance and color, we adopt a different approach since we can optimize a consistent set of Gaussian attributes during this phase. Specifically, we sample the spherical harmonics (SHs) from the appearance model $\mathcal{A}$ and make them a learnable, per-Gaussian variable as in the original 3DGS. Moreover, the Gaussian covariance is parameterized into learnable 2D scaling $\hat{\boldsymbol{s}}\in\mathbb{R}^{2}$ and a complex number $\boldsymbol{a}i+\boldsymbol{b}$ representing rotation, which are similar to the process in SuGaR [7]. It should be noted that this process is optional, as in many cases, the quality of the learned geometry and appearance without refinement is already satisfactory.
5 Experiments
In this section, we first detail the implementation across different stages and describe the datasets used for evaluation.
To assess the effectiveness of our method, we compare it with models that specialize in novel view synthesis and surface reconstruction tasks, respectively. Note that the mesh is not used in the rendering of our method. Instead, it serves more as a constraint, ensuring that the Gaussians are aligned with the mesh to faithfully represent the geometry while also reproducing the scene’s appearance. We will demonstrate that, even with this constraint, our model can approximate or even surpass the performance of models without these constraints.
Furthermore, we showcase potential applications in object manipulation, and then discuss the adaptability to scene updates and ablation studies in the method analysis section.
5.1 Implementation details
The proposed method is implemented using PyTorch and the rasterization toolkit from 3DGS [5]. Experiments are conducted on an RTX3090 GPU with 24GB memory.
5.1.1 Initialization
We train the original 3DGS model for 20,000 iterations to obtain a set of Gaussians, whose center positions are treated as a point cloud. We then employ Alpha Shapes [34] to extract a coarse mesh and make it watertight [35], which is subsequently used to construct an SDF grid. Additionally, the background Gaussians are retained and continuously trained in the later stages.
Our approach differs from others [12, 13] that randomly initialize the SDF, which would cause an OOM issue when the scene becomes large. Thanks to the fast optimization and the discrete nature of 3DGS, we are able to capture the coarse geometry using only 5-20 minutes.
5.1.2 Joint Learning of Mesh and Gaussians
The resolution of the used SDF grid depends on the scene. For Mip-NeRF360 scenes, we use $150^{3}$ to $250^{3}$ grid cells for the foreground, attaching 3 Gaussians to every face. For NeRF-Synthetic scenes, we use $100^{3}$ grid cells and attach 3 Gaussians to every face. For the appearance model, we adopt an MLP with 2 hidden layers, each comprising 32 neurons. This stage requires 10k iterations for optimization, totally using 15-30 minutes.
5.1.3 Refinement
Before the refinement stage, we perform subdivision and decimation on the mesh to unify the number of faces (0.1M for single objects, 0.5M for real-world scenes). We attach 6 Gaussians to each face and compute their covariances and colors using the adaptive Gaussian construction and appearance model, with minimal impact on model performance. The refinement process consists of 5-10k iterations and requires 10-20 minutes.
5.2 Datasets and Metrics
The NeRF-Synthetic dataset [4] includes 8 scenes with 360° viewpoint settings of images. Besides testing novel view synthesis on this dataset, we quantitatively evaluate reconstructed surfaces using the provided ground-truth mesh. We also evaluate our method on 7 of the 9 real-world scenes in the Mip-NeRF360 dataset [36], excluding Flowers and Treehill due to licensing restrictions. The strategy for splitting subsets for training and testing follows that of 3DGS [5].
For novel view synthesis, we employ the standard PSNR, L-PIPS, and SSIM metrics. In the surface reconstruction task, we compute the Chamfer Distance (CD) between the extracted mesh and the ground truth.
Table 1: Per-scene quantitative comparisons on NeRF-Synthetic dataset [4].
Method Chair Drums Ficus Hotdog
[…]
without mesh
NeRF [4] 33.00 0.967 0.046 25.01 0.925 0.091 30.13 0.964 0.044 36.18 0.974 0.121
Plenoxels [16] 33.98 0.977 0.031 25.35 0.933 0.067 31.83 0.976 0.026 36.43 0.980 0.037
3DGS [5] 35.82 0.987 0.012 26.17 0.954 0.037 34.83 0.987 0.012 37.67 0.985 0.020
with mesh
NVdiffRec [12] 31.60 0.969 0.034 24.10 0.916 0.065 30.88 0.970 0.041 33.04 0.973 0.033
NeuManifold [33] 34.39 0.981 0.014 25.39 0.939 0.072 31.91 0.978 0.028 35.69 0.979 0.036
NeRF2Mesh [21] 34.25 0.978 0.031 25.04 0.926 0.084 30.08 0.967 0.046 35.70 0.974 0.058
SuGaR-15K [7] 35.13 0.983 0.014 25.44 0.943 0.057 32.75 0.982 0.018 36.67 0.980 0.021
Ours 35.40 0.986 0.013 25.75 0.952 0.041 34.01 0.986 0.014 36.80 0.985 0.018
Method Lego Materials Mic Ship
[…]
without mesh
NeRF [4] 32.54 0.961 0.050 29.62 0.949 0.063 32.91 0.980 0.028 28.65 0.856 0.206
Plenoxels [16] 34.10 0.975 0.028 29.14 0.949 0.057 33.26 0.985 0.015 29.62 0.890 0.134
3DGS [5] 35.69 0.983 0.016 30.00 0.960 0.034 35.34 0.991 0.006 30.87 0.907 0.106
with mesh
NVdiffRec [12] 29.14 0.949 0.042 26.74 0.923 0.060 30.78 0.977 0.024 26.12 0.833 0.080
NeuManifold [33] 34.00 0.977 0.024 26.69 0.924 0.115 33.40 0.986 0.012 28.63 0.875 0.168
NeRF2Mesh [21] 34.90 0.977 0.025 26.26 0.906 0.111 32.63 0.979 0.038 29.47 0.875 0.138
SuGaR-15K [7] 34.87 0.981 0.015 27.86 0.944 0.046 34.81 0.989 0.008 29.22 0.880 0.103
Ours 34.94 0.982 0.016 27.93 0.950 0.043 34.89 0.991 0.007 28.32 0.883 0.121
Table 2: Quantitative comparisons on Mip-NeRF360 dataset [36].
Method Indoor scenes Outdoor scenes All scenes
[…]
without mesh
Plenoxels [16] 24.83 0.766 0.426 22.02 0.542 0.465 23.62 0.670 0.443
INGP-Base [18] 28.65 0.840 0.281 23.47 0.571 0.416 26.43 0.725 0.339
3DGS [5] 30.41 0.920 0.189 26.40 0.805 0.173 28.69 0.870 0.182
with mesh
NeRFMeshing [29] 23.83 - - 22.23 - - 23.15 - -
BakedSDF [2] 27.06 0.836 0.258 - - - - - -
NeuManifold [33] 27.16 0.813 0.316 - - - - - -
NeRF2Mesh [21] - - - 22.74 0.523 0.457 - - -
SuGaR-15K [7] 29.43 0.910 0.216 24.40 0.699 0.301 27.27 0.820 0.253
Ours 29.33 0.910 0.193 25.17 0.754 0.237 27.54 0.843 0.212
Figure 5: Qualitative comparisons with baseline methods (SuGaR [7], NeRF2Mesh [21]) on the NeRF-Synthetic [4] and Mip-NeRF360 dataset [36].
5.3 Novel View Synthesis
We first evaluate our method on the novel view synthesis task. Tabs. 1 and 2 present the quantitative results on NeRF-Synthetic and Mip-NeRF360 datasets. To ensure a fair comparison, we employed a white background for the NeRF-Synthetic dataset. Our approach is compared against various recent methods, including those with reconstructed mesh and those without.
For the NeRF-Synthetic dataset, our method achieves state-of-the-art rendering quality among methods reconstructing a mesh, and even outperforms some models that solely focus on rendering quality. Furthermore, there is only a slight decrease in performance compared to the original 3DGS [5]. For the Mip-NeRF360 dataset, our method is comparable to or surpasses the state-of-the-art methods [7, 21]. Fig. 5 includes qualitative comparisons between images rendered by our method and the baselines, demonstrating that our method captures better scene details.
Figure 6: Qualitative comparisons of reconstructed meshes with SuGaR [7] on the NeRF-Synthetic [4] and Mip-NeRF360 [36] datasets.
Table 3: Quantitative Evaluation of Mesh Quality and Efficiency on the NeRF-Synthetic [4] dataset.
Method CD ($10^{-3}$) #Faces PSNR Train Time
NVdiffRec [12] 8.15 80k 29.05 52 mins
NeRF2Mesh [21] 5.06 192k 31.04 46 mins
SuGaR-15K [7] 8.62 1000k 32.09 103 mins
Ours (w/o refine) 8.50 141k 27.68 16 mins
Ours 7.27 100k 32.26 28 mins
‘CD’: Chamfer Distance between the ground truth and reconstructed meshes.
5.4 Surface Reconstruction
Although our focus is on efficiently learning a hybrid representation that ensures alignment between Gaussians and mesh faces, the surfaces reconstructed by our method are of high quality. In Fig. 6, we visualize the reconstructed meshes and their normal maps, comparing our method against the baseline. The results demonstrate that our method effectively captures the surface details of a scene, whereas SuGaR [7] may produce floating patches and holes in the mesh.
For a quantitative evaluation of the meshes, we compare the reconstructed mesh to the ground truth from the NeRF-Synthetic dataset using the Chamfer Distance (CD) metric. The computation method follows NeRF2Mesh [21], where 2.5M points are sampled from the surfaces of both the ground truth and the reconstructed object through ray casting. For a more comprehensive assessment, Tab. 3 presents the CD results, the number of faces, the rendering PSNR, and the training time. The results demonstrate that our reconstructed meshes achieve both high quality and training efficiency. Additionally, our Gaussian-based appearance model enables high-quality rendering even with fewer faces compared to NeRF2Mesh.
Furthermore, the learned mesh without refinement is already of satisfactory quality. This indicates that for applications requiring only an untextured mesh, our method can produce excellent results very efficiently, with a training time of just 16 minutes.
Figure 7: Object deformation by hybrid representation of mesh and Gaussians. We employ four Blender modifiers to manipulate the mesh, subsequently render perspectives using the bound Gaussians.
5.5 Object Deformation by Hybrid Representation
As discussed in our introduction section, the hybrid representation not only benefits from the high-quality and efficient rendering of Gaussians but also offers the convenience of manipulation through the mesh. To demonstrate the potential applications of this work, we selected object deformation tasks for testing. Specifically, we deform the learned mesh using four common mesh modifiers (Twist, Stretch, Bend, and Taper) in Blender and correspondingly adjust the bound Gaussians using the technique described in Sec. 4.2. We then use the Gaussian-based rasterization to render images of the deformed object, as shown in Fig. 7. More advanced operations, as explored in [9, 37], are also compatible with the representation we have obtained.
5.6 Method Analysis
Figure 8: Adaptation to scene updates. (a) Models initially trained on the original scene. (b) and (c): Models adapted to the updated scene for 500 and 5000 steps, respectively. (d): The meshes learned on the updated scene.
Table 4: Evaluation on Efficiency. We report the training time and PSNR metric across different methods.
Method NeRF-Synthetic [4] Mip-NeRF360 [36]
[…]
NVdiffRec [12] 29.05 52 mins - -
NeRF2Mesh [21] 31.04 46 mins - -
SuGaR-15K [7] 32.09 103 mins 27.27 132 mins
Ours 32.26 28 mins 27.54 61 mins
‘-’ indicates that the method does not have results reported in their paper and fails to reconstruct all scenes of the dataset using their published codes.
Table 5: Ablation studies on adaptive covariance and refinement stage. We report the average PSNR across all scenes of the NeRF-Synthetic [4] and Mip-NeRF360 [36] datasets.
Adaptive Cov. Refine NeRF-Synthetic Mip-NeRF360
x $\checkmark$ 32.02 27.19
$\checkmark$ x 27.78 24.44
$\checkmark$ $\checkmark$ 32.26 27.54
5.6.1 Efficiency
We assess our method’s efficiency by comparing the average time taken to reconstruct scenes from the NeRF-Synthetic and Mip-NeRF360 datasets. Both our method and the baseline methods are tested under identical hardware conditions. The results in Tab. 4 reveal that our method is not only faster but also provides superior rendering quality.
5.6.2 Adaptability in Scene Updates
One unique feature of our method is the simultaneous learning of topology (defined by mesh faces) and Gaussians. In contrast, the existing method, SuGaR [7, 15, 37], separates the learning of topology and Gaussians, as the mesh extraction process (Poisson Surface Reconstruction) is non-differentiable for Gaussian rasterization. In practice, when part of the scene requires updating, our model can directly adapt to these changes without re-learning from scratch, whereas the existing method cannot. Fig. 8 presents the results of handling scene updates using our method and SuGaR. Our method successfully updates the geometry and appearances of the modified regions, while SuGaR struggles to adapt to new scene geometries with its fixed topology.
5.6.3 Ablation Studies
In this section, we examine the impact of three factors: 1) the adaptive Gaussian covariance, 2) the refinement stage, and 3) the number of Gaussians per face. To evaluate the effect of adaptive Gaussian covariance, we set $\boldsymbol{M}$ in Eq. (10) to be an identity matrix. Tab. 5 shows that both adaptive Gaussian covariance and the refinement stage significantly enhance rendering quality.
The choice of the number of Gaussians per face involves a trade-off between computational consumption and performance. As shown in Tab. 6, increasing the number of Gaussians per face benefits rendering quality. To control the memory footprint and training time within an acceptable range while still achieving excellent rendering quality, we choose 3 Gaussians per face during the joint learning of mesh and Gaussian appearance, but use 6 Gaussians for the refinement stage.
Table 6: Ablation Study on the Number of Gaussians per Face. This table shows the PSNR results on the Mip-NeRF360 [36] dataset.
6 Conclusion
We have proposed a novel learning method to capture comprehensive 3D scene information from multiple views. The method simultaneously extracts both the geometry and the physical properties affecting the observed appearance. The geometry is extracted in the explicit form of triangle meshes. The appearance properties are encoded in 3D Gaussians that are bound to the mesh faces. Thanks to the 3DGS-based differentiable rendering, we are able to establish an effective and efficient learning procedure by directly optimizing the photometric loss. Experiments verify that the resulting representation enjoys both high-quality rendering and controllability.
[…]
Data availability
The datasets analyzed during the current study are available in these repositories: NeRF-Synthetic https://github.com/bmild/nerf; Mip-NeRF360 https://jonbarron.info/mipnerf360.
Appendix A
Table 7: Per-scene quantitative results on Mip-NeRF360 dataset [36].
Metric Bicycle Bonsai Counter Garden Kitchen Room Stump Mean
PSNR 24.18 31.47 27.16 26.41 28.96 29.71 24.91 27.54
SSIM 0.722 0.945 0.888 0.829 0.898 0.908 0.712 0.843
LPIPS 0.260 0.175 0.212 0.161 0.162 0.221 0.291 0.212
A.1 Details for SDF Grid Optimization
This section includes the methods and configurations used to mitigate the storage and computational demands of the SDF grid:
[…]
Visible Node Optimization For the given grid nodes $\{\boldsymbol{x_{n}}\}_{n=1}^{N}$, the visible subset is calculated by:
$\{\boldsymbol{\hat{x}_{m}}\}:=\left\{\boldsymbol{x_{n}}\big{|}\text{NDC}(\boldsymbol{x_{n}})\in[-1,1]^{3}\right\},$ (12)
where $\text{NDC}(\boldsymbol{x_{n}})$ is the transformation from world-space coordinates $\boldsymbol{x_{n}}$ to their normalized device coordinates (NDC) via the projection matrix. The marching algorithm is only applied to cells whose eight corner nodes are within $\{\boldsymbol{\hat{x}_{m}}\}$.
[…]
Coarse-to-fine Grid Given the bounding box (BBox) with size ($L_{x}$, $L_{y}$, $L_{z}$) that tightly encloses the foreground region $U$, and an expected cell count $C$, the cubic cell size of the SDF grid is set to $(L_{x}/s)\times(L_{y}/s)\times(L_{z}/s)$, where $s=\sqrt[3]{L_{x}\cdot L_{y}\cdot L_{z}/C}$. During optimization, we perform 4 coarse-to-fine operations, incrementally increasing the cell count by 1.5×1.5$\times$1.5 × each step until the target count $C$ is reached.
A.2 Configurations of Face Gaussians
As discussed in Sec. 4.2, each face is associated with $K$ Gaussians. The choice of $K$ influences the barycentric coordinates of the Gaussian centers and the radius $r$ employed in the covariance matrix $\boldsymbol{\Sigma}_{e}$ in Eq. (10). For $K=3$, we use
$\boldsymbol{\xi}_{1}$ $=\left[(3-\sqrt{3})/6,(3-\sqrt{3})/6,\sqrt{3}/3\right],$
$\boldsymbol{\xi}_{2}$ $=\left[(3-\sqrt{3})/6,\sqrt{3}/3,(3-\sqrt{3})/6\right],$
$\boldsymbol{\xi}_{3}$ $=\left[\sqrt{3}/3,(3-\sqrt{3})/6,(3-\sqrt{3})/6\right],$
and radius $r=\|\boldsymbol{v}_{1}-\boldsymbol{v}_{2}\|/(2\sqrt{3}+2)$. For $K=6$, we use
$\boldsymbol{\xi}_{1}$ $=\left[(3-2\sqrt{3})/6,(3-2\sqrt{3})/6,2\sqrt{3}/3\right],$
$\boldsymbol{\xi}_{2}$ $=\left[(3+2\sqrt{3})/12,(3-\sqrt{3})/6,(3+2\sqrt{3})/12\right],$
$\boldsymbol{\xi}_{3}$ $=\left[(3-\sqrt{3})/6,(3+2\sqrt{3})/12,(3+2\sqrt{3})/12\right],$
$\boldsymbol{\xi}_{4}$ $=\left[2\sqrt{3}/3,(3-2\sqrt{3})/6,(3-2\sqrt{3})/6\right],$
$\boldsymbol{\xi}_{5}$ $=\left[(3+2\sqrt{3})/12,(3+2\sqrt{3})/12,(3-\sqrt{3})/6\right],$
$\boldsymbol{\xi}_{6}$ $=\left[(3-2\sqrt{3})/6,2\sqrt{3}/3,(3-2\sqrt{3})/6\right],$
and radius $r=\|\boldsymbol{v}_{1}-\boldsymbol{v}_{2}\|/(2\sqrt{3}+4)$.
A.3 Derivation of Adaptive Gaussian Covariance
Figure 9: The linear transformation applied on Gaussians for a irregular triangle.
This section presents how to obtain the matrix $\boldsymbol{M}$ used in (10) for a target triangle $(\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3})$.
As shown in Fig. 9, the linear transformation applied on a reference equilateral triangle will result in the target triangle. In the local coordinate system defined in (9), the target triangle is denoted as $\boldsymbol{T}=[\boldsymbol{v}^{\prime}_{1},\boldsymbol{v}^{\prime}_{2},\boldsymbol{v}^{\prime}_{3}]$, and the reference equilateral triangle as $\boldsymbol{E}=[\boldsymbol{v}^{\prime}_{1},\boldsymbol{v}^{\prime}_{2},\hat{\boldsymbol{v}}^{\prime}_{3}]$, where $\hat{\boldsymbol{v}}^{\prime}_{3}=l\cdot[0,\frac{1}{2},\frac{\sqrt{3}}{2}]^{T}$, with $l=\|\boldsymbol{v}_{2}-\boldsymbol{v}_{1}\|$.
By solving the equation $\boldsymbol{M}\boldsymbol{E}=\boldsymbol{T}$, we obtain:
$\boldsymbol{M}=\begin{bmatrix}1&0&0\\ 0&1&\frac{2\boldsymbol{v}^{\prime(2)}_{3}-l}{\sqrt{3}l}\\ 0&0&\frac{2\boldsymbol{v}^{\prime(3)}_{3}}{\sqrt{3}l}\\ \end{bmatrix},$ (13)
where $\boldsymbol{v}^{\prime(i)}_{3}$ denotes the $i$-th component of $\boldsymbol{v}^{\prime}_{3}$.
A.4 More Results
Tab. 7 includes the per-scene results on Mip-NeRF360 dataset [36].